OpenStudy (juan1857):

What are the expressions in factored form? 1)x^2-64 2)9x^2-18x+9 3)16x^2-25

jhonyy9 (jhonyy9):

in case of first just yous formula of difference of two squares - do you know it ?

jhonyy9 (jhonyy9):

ohhh and for 3rd same too

OpenStudy (juan1857):

no, is there a broad method to find out the answers, and is it the easiest

jhonyy9 (jhonyy9):

so than you know this formuela this is the easy way sure this is the formula for difference of two squares a^2 -b^2 =(a-b)(a+b)

OpenStudy (juan1857):

looks familiar

jhonyy9 (jhonyy9):

just use it for the first will be x^2 -64 = (x-8((x+8) yes ?

jhonyy9 (jhonyy9):

for 3rd will be the same easy 16x^2 -25 = (4x)^2 -5^2 = ?

OpenStudy (juan1857):

so do you square root always?

jhonyy9 (jhonyy9):

sorry above changed the paranthese there is (x-8)(x+8)

OpenStudy (juan1857):

oh ok

jhonyy9 (jhonyy9):

in this case yes bc. there are square perfect and this is why you can using this formula easy

OpenStudy (juan1857):

what if they don't square out perfectly?

jhonyy9 (jhonyy9):

so for second one 9x^2-18x+9 =0 how you see what can you factorizing out from every 3 terms of quadratic ?

OpenStudy (juan1857):

a 9

jhonyy9 (jhonyy9):

yes please make it and see what will get - ok. ?

OpenStudy (juan1857):

ok

jhonyy9 (jhonyy9):

here please

OpenStudy (juan1857):

oh well I ended up with 9(x^2-2x+1) I don't know what to do after

jhonyy9 (jhonyy9):

x^2 -2x +1 =0 so you dont see it nothing here now just bc. this quadratic is a square perfect or not ?

jhonyy9 (jhonyy9):

looke please (x-1)^2 = ? how many will give you ?

OpenStudy (juan1857):

oh yeah, but what do you do with the 9?

jhonyy9 (jhonyy9):

so 9 will be there in front of the first paranthese - ok ?

OpenStudy (juan1857):

ok so would the answer be (9x-1)(x-1)??

jhonyy9 (jhonyy9):

no i said above in front of the first paranthese not inside

OpenStudy (juan1857):

so 9(x-1)(x-1)?

jhonyy9 (jhonyy9):

yes but hence is more nicely 9(x-1)^2 or not ?

OpenStudy (juan1857):

yeah

jhonyy9 (jhonyy9):

ok was my pleasure good luck bye bye

OpenStudy (juan1857):

thank you

jhonyy9 (jhonyy9):

np anytime